What is the potential at a distance of 0.0529 nm from a proton?

a.13.6 no

b.-13.6 no

c.27.2 V

d. -27.2 no

e.9.11 pC
Answer #1

The electric potential from a distance r for a charge q is given by the equation is

v = q/(4*pi*\epsilon *r)

Here q = 1.6*10^{-19}   and \epsilon = 8.85*10^{-12} and r = 0.0529 nm

By substituting the values in the equation we will get

v = 27.2 V